OpenStudy (anonymous):

"for a given phase shift the higher is the frequency, the smaller is the required time shift"...please explain this statement..is this in any indirect way related to doppler's law?

OpenStudy (vincent-lyon.fr):

Interferences? phase shift \(\Delta \phi\) is to \(2\pi\) what time shift \(\Delta \tau\) is to time period \(T\) Hence \(\Delta \tau=\Delta \phi\times T/2\pi=\Delta \phi /2\pi\nu\) You can see that for a given \(\Delta \phi\) , \(\Delta \tau\) is inversely proportional to frequency \(\nu\).